OpenStudy (anonymous):

The regular price for a printer is $200. After discount it is $120. What is the discount do the person get?

OpenStudy (anonymous):

how much did the guy save?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

er, how many dollars did the guy save?

OpenStudy (anonymous):

im asking lol, this is how you find out the discount.

OpenStudy (anonymous):

No what is the discount the guy get.

OpenStudy (anonymous):

200-120=?

OpenStudy (anonymous):

No percent of the discount.

OpenStudy (anonymous):

\[(x/100)*200=80\]

OpenStudy (anonymous):

Answer the give u..... 4%, 50%,40%,75%

OpenStudy (anonymous):

40%

OpenStudy (anonymous):

\[1-\frac{120}{200}=0.4 =40\%\]

OpenStudy (anonymous):

Thanks.